# FreeGAD: A Training-Free yet Effective Approach for Graph Anomaly Detection

### v1

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11ccs: Computing methodologies Neural networksccs: Mathematics of computing Graph algorithms
@@ 2.2. Graph Neural Networks
Graph neural networks (GNNs) are a neural network architecture designed for graph-structured data, employing a message-passing mechanism with two key operations: propagation (P), which facilitates information exchange between neighboring nodes, and transformation (T), which updates node embeddings through learnable non-linear projections (Wu et al., 2020). Due to their powerful capabilities, GNNs have been successfully applied across a wide range of tasks (Li et al., 2024; Liu et al., 2024c, a; Shen et al., 2025; Li et al., 2025). Different GNN models leverage diverse propagations and aggregation functions. Frequently utilized P functions comprise averaging aggregation (Kipf and Welling, 2017), attention aggregation (Veličković et al., 2017), and LSTM aggregation (Hamilton et al., 2017), while T operations are typically implemented as one or more perceptron layers followed by non-linear activations.
@@ 4.1. Affinity-Gated Residual Encoder
In the first step of FreeGAD, we design an encoder to generate representations of nodes. The representations are expected to incorporate the semantic and structural information of each node, providing anomaly-related cues for effective abnormality prediction. The state-of-the-art GAD methods (e.g., CoLA (Liu et al., 2021) and TAM (Qiao and Pang, 2023)) usually employ GCN (Kipf and Welling, 2017) as encoder, which may suffer from two limitations. Firstly, stacking many layers can degrade the representation capability of GCN due to over-smoothing (Chen et al., 2020) and model degradation issues (Zhang et al., 2022). This limitation hinders GCN from capturing high-order information from long-range nodes, which is crucial for detecting real-world anomalies. Secondly, GCN learns representations by simply averaging ego and neighbor information, which may fail to capture anomaly-sensitive knowledge (e.g. affinity (Qiao and Pang, 2023)).
@@ 5. Experiments
Metric Method DataSet
Amazon Reddit Tolokers YelpChi T-Finance Questions Elliptic Cora BlogCatalog Flickr

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761125isbn: 979-8-4007-2040-6/2025/11ccs: Computing methodologies Neural networksccs: Mathematics of computing Graph algorithms
@@ 2.2. Graph Neural Networks
Graph neural networks (GNNs) are a neural network architecture designed for graph-structured data, employing a message-passing mechanism with two key operations: propagation (P), which facilitates information exchange between neighboring nodes, and transformation (T), which updates node embeddings through learnable non-linear projections (Wu et al., 2020). Due to their powerful capabilities, GNNs have been successfully applied across a wide range of tasks (Li et al., 2024; Liu et al., 2024c, a; Shen et al., 2025; Li et al., 2025). Different GNN models leverage diverse propagations and aggregation functions. Frequently utilized P functions comprise averaging aggregation (Kipf, 2016), attention aggregation (Veličković et al., 2017), and LSTM aggregation (Hamilton et al., 2017), while T operations are typically implemented as one or more perceptron layers followed by non-linear activations.
@@ 4.1. Affinity-Gated Residual Encoder
In the first step of FreeGAD, we design an encoder to generate representations of nodes. The representations are expected to incorporate the semantic and structural information of each node, providing anomaly-related cues for effective abnormality prediction. The state-of-the-art GAD methods (e.g., CoLA (Liu et al., 2021) and TAM (Qiao and Pang, 2023)) usually employ GCN (Kipf, 2016) as encoder, which may suffer from two limitations. Firstly, stacking many layers can degrade the representation capability of GCN due to over-smoothing (Chen et al., 2020) and model degradation issues (Zhang et al., 2022). This limitation hinders GCN from capturing high-order information from long-range nodes, which is crucial for detecting real-world anomalies. Secondly, GCN learns representations by simply averaging ego and neighbor information, which may fail to capture anomaly-sensitive knowledge (e.g. affinity (Qiao and Pang, 2023)).
@@ 5. Experiments
Metric Method Amazon Reddit Tolokers YelpChi T-Finance Questions Elliptic Cora BlogCatalog Flickr